# Out-of-Distribution Generalization in the ARC-AGI Domain: Comparing Execution-Guided Neural Program Synthesis and Test-Time Fine-Tuning

### v1

@@ Introduction
While ARC-AGI may not capture all aspects of AGI, it offers a concrete test of a critical capability: solving structurally novel reasoning tasks by recombining known primitives. This could be one of the main bottlenecks preventing us from reaching AGI in any deep or strong sense of the word. Indeed, cutting edge AI systems fail to achieve results above 20% on ARC-AGI-2 at the time of this writing. The fact that humans solve these problems while our best AI systems do not is direct, practical evidence of a fundamental missing functionality with respect to AGI.
@@
    Given the selected node, the target grids, and the trained model: generate all non-zero probability token sequences for the next instruction step.
@@
Each instruction is formatted as a token sequence (see Table 1):
@@
Table 2: Token ranges in the DSL vocabulary. The vocabulary includes integer constants, DSL primitives, grid attribute selectors, and references to previously defined state variables.
#1. [equal, <SEP>, N+0, ’.c’, <ARGSEP>, 0, <EOS>]     $\rightarrow N+1$
#2. [switch, <SEP>, N+1, <ARGSEP>, 0, <ARGSEP>, 2, <EOS>]     $\rightarrow N+2$
#3. [del, <SEP>, N+1, <EOS>]     $\rightarrow N+3$
#4. [set_pixels, <SEP>, N+0, <ARGSEP>, N+0.’.x’, <ARGSEP>, N+0.’.y’, <ARGSEP>, N+2, <EOS>]
@@
The example program in Table 3 contains four instruction steps. Its effect is to change all non-zero pixels of the grid to the color $2$. The first instruction returns a boolean list for each pixel in the input grid set (reference: $N+0$), set to true if 00, false otherwise. Instruction $2$ is an ”if/else” statement going over each element of $N+1$. When its value is True, it outputs 00, otherwise it outputs $2$. The third instruction removes from memory the now obsolete $N+1$. The last instruction sets all pixels of the input grid set $(N+0)$, at positions determined by .x and .y attributes of $N+0$ (i.e., all pixels are modified), to the color value specified in what was previously $N+2$, but is now $N+1$ due to the previous del operation.
@@
The training data for each task is generated by randomly sampling subgrids of ARC-AGI training tasks, and by applying the task’s ground truth program to the inputs in order to obtain the target grids. The program synthesis models are trained on $200,000$ task examples each containing exactly one input/output pair, each example designed to be non-trivial and not underdetermined. The task is considered solved if it finds a program that generates the output from the input. Only one run per task example was done. The DSL for the experiments #1 and #2 contain $34$ primitives, while the experiment #3 used a DSL of $45$ primitives.
The TTFT model was pretrained on a larger amount of pair examples: $100,000$ task instances each having 8 input/output demonstration pairs, for a total of $800,000$ distinct pair examples. TTFT itself is done on task instances with 3 pairs. In that case, because it is a transductive approach, we present it with a test input grid, and it must generate the corresponding output according to the task’s logic. The voting procedure is done over 8 different inference attempts.
@@
Total 10% 42.86% 80% 10% 10%
Table 4: Success rate over 10 samples for each OOD task and each algorithm, with a 3-minute budget per task sample.
@@ Acknowledgments
3:  $start\_time=time()$
@@
5:     $output=execute\_program(selected)$
@@
12:        $log\_prob=calculate\_joint\_prob(prog)$
@@
15:     $next\_best=$ Dequeue program queue
16:     $leaf=$ program leaf node for $next\_best$
@@
In the above pseudocode, $SearchTreeNode$ is the class representing a node in the tree structure. The first argument is the state variable if we already have it, and the second argument is the parent node. The node’s state variable can be set to None and later updated from inside the $execute\_program$ call. The latter executes the instruction step associated with the node, gets the resulting output, and updates the node’s internal state variable. It also returns the result to verify if the goal was reached.
The $enumerate$ function aggregates encoder output from the parent node states and the target grid $Y$, and then uses the model $M$ to recursively query token probabilities and generate all non-zero probability token sequences for the next instruction step. It returns these new generated programs. Suppose $enumerate$ generates $42$ different possible instruction steps at this stage in the program, this means that we add $42$ new programs to the program queue: each starting with the same instruction steps for the node parents trajectory, and ending with the newly generated instruction steps.

### v2

@@ Introduction
While ARC-AGI may not capture all aspects of AGI, it offers a concrete test of a critical capability: solving structurally novel reasoning tasks by recombining known primitives. This could be one of the main bottlenecks preventing us from reaching AGI in any deep or strong sense of the word. Indeed, cutting edge AI systems fail to achieve results above 29% on ARC-AGI-2 at the time of this writing. The fact that humans solve these problems while our best AI systems do not is direct, practical evidence of a fundamental missing functionality with respect to AGI.
@@
    Given the selected node, the target grids, and the trained model: generate all token sequences for the next instruction step (whose probability >> some threshold, 0.01 in our case).
@@
Table 2: Token ranges in the DSL vocabulary. N is the total number of primitives (including the constants and object attributes) in the DSL. The vocabulary includes integer constants, DSL primitives, object attributes, and references to previously defined state variables.
The example program in Table 3 contains four instruction steps. Its effect is to change all non-zero pixels of the grid to the color $2$. The first instruction returns a boolean list for each pixel in the input grid set (reference: $N+0$), set to true if 0, false otherwise. Instruction $2$ is an ”if/else” statement going over each element of $N+1$. When its value is True, it outputs 0, otherwise it outputs $2$. The third instruction removes from memory the now obsolete $N+1$. The last instruction sets all pixels of the input grid set $(N+0)$, at positions determined by .x and .y attributes of $N+0$ (i.e., all pixels are modified), to the color value specified in what was previously $N+2$, but is now $N+1$ due to the previous del operation.
[…]
#3. [del, <SEP>, N+1, <EOS>]
#4. [set_pixels, <SEP>, N+0, <ARGSEP>, N+0, <.x>, <ARGSEP>, N+0, <.y>, <ARGSEP>, N+1, <EOS>]
@@
The training data for each task is generated by randomly sampling subgrids of ARC-AGI training tasks, and by applying the task’s ground truth program to the inputs in order to obtain the target grids. The program synthesis models are trained on $200,000$ task examples each containing exactly one input/output pair, each example designed to be non-trivial and not underdetermined. The task is considered solved if it finds a program that generates the outputs (including a test output) from the input. Five runs per task example were done to evaluate variance in results (min/max/median). The DSL for the experiments #1 and #2 contain $34$ primitives, while the experiment #3 used a DSL of $45$ primitives.
The TTFT model was pretrained on a larger amount of pair examples: $100,000$ task instances each having 8 input/output demonstration pairs, for a total of $800,000$ distinct pair examples. TTFT itself is done on task instances with 3 pairs. The voting procedure, when evaluating the test output, is done over 8 different inference attempts. The most frequent output is kept.
The comparison between TTFT and GridCoder2 is fair because they both have 3 minutes of test-time search (or gradient descent in the case of TTFT). If anything, the experiment is unfair to GridCoder2 because it is trained on only $200,000$ whereas TTFT is trained on $800,000$ tasks. The 3 minutes time budget was selected to match the average task solution time allowed per task on the ARC-AGI Kaggle competition, and is therefore not arbitrary, nor was it selected to favor GridCoder2. Qwen2-0.5B-Instruct is the LLM used by Omni-ARC to reach its performance in the 2024 competition, so it has more than enough capacity to learn what it needs to learn for these experiments.
@@
Min 10% 40.06% 78.4% 9.8% 14.29%
Max 10% 43% 82.8% 12.3% 14.29%
Median 10% 42.86% 80% 10% 14.29%
Table 4: Median success rate over 10 samples for each OOD task and each algorithm, with a 3-minute budget per task sample. Total min, max and median over 5 distinct runs is shown as well.
@@ Acknowledgments
3:start_time=time()start\_time=time()
@@
5:  output=execute_program(selected)output=execute\_program(selected)
@@
12:   log_prob=calculate_joint_prob(prog)log\_prob=calculate\_joint\_prob(prog)
@@
15:  next_best=next\_best= Dequeue program queue
16:  $leaf=$ program leaf node for next_bestnext\_best
@@
In the pseudocode, $SearchTreeNode$ is the class representing a node in the tree structure. The first argument is the state variable if we already have it, and the second argument is the parent node. The node’s state variable can be set to None and later updated from inside the execute_programexecute\_program call. The latter executes the instruction step associated with the node, gets the resulting output, and updates the node’s internal state variable. It also returns the result to verify if the goal was reached.
The enumerateenumerate function aggregates encoder output from the parent node states and the target grid $Y$, and then uses the model $M$ to recursively query token probabilities and return possible token sequences for the next instruction step whose probability is greater than some threshold. We use 0.01 in our experiments. Also, if the number of generated possible sequences exceeds a threshold k, we only keep the top k (k = 1M in our experiments). Finally, there is a maximum token sequence length of 40 tokens.
Suppose enumerateenumerate generates $42$ different possible instruction steps at this stage in the program, this means that we add $42$ new programs to the program queue: each starting with the same instruction steps for the node parents trajectory, and ending with the newly generated instruction steps.
Not presented in the pseudo-code (for simplicity) are further search constraints such as maximum tree depth (20 in our experiments) and a memory limit of 5. This memory limit refers to the maximum number of intermediate state variables that can be kept in the program’s working memory at all times. When that memory limit is reached, only the del primitive is allowed to clear up some memory.